# Hidden ferromagnetism of centrosymmetric antiferromagnets

### v1

@@ Acknowledgement
Appendix A Elimination of weak ferromagnetic components of the spin-orbit coupling
@@
Then, let us consider the unitary transformation of the sublattice $2$, $\hat{U}=e^{i\varphi\boldsymbol{n}\cdot\hat{\boldsymbol{\sigma}}}=\cos\frac{\varphi}{2}\hat{\mathbb{1}}+i\boldsymbol{n}\cdot\hat{\boldsymbol{\sigma}}\sin\frac{\varphi}{2}$, such that $\left(t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}\right)\hat{U}\approx\tilde{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+it^{y}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\sigma}_{y}$. In principle, by using such transformation, one can completely eliminate the SO term $i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ separately in each bond [29, 50]. However, this cannot be done simultaneously for all the bonds, unless the parameters of the SO coupling are the same in all these bonds, as the $x$ and $z$ components of the SO coupling responsible for the weak ferromagnetism, for which $t^{x}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv t_{x}$ and $t^{z}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv t_{z}$.
Thus, assuming $\boldsymbol{n}=\frac{(t_{x},0,t_{z})}{\sqrt{t_{x}^{2}+t_{z}^{2}}}$, which does not depend on the alternating $y$-component, one can find the following equations for $\varphi$:
@@ Appendix B Electronic structure and minimal model for VF4 and CuF2
The transfer integrals operating within the same sublattice were additionally averaged over two sublattices. Namely, using the transfer integrals in the sublattices $I$ and $II$, one can define $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{1}{2}(t^{I}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}+t^{II}_{\boldsymbol{R},\boldsymbol{R}^{\prime}})$ (the regular contribution) and $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{1}{2}(t^{I}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}-t^{II}_{\boldsymbol{R},\boldsymbol{R}^{\prime}})$ (the altermagnetic contribution). The approximation consists in neglecting $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$. This artificially suppress the splitting of the AFM bands, but allows us to employ the generalized Bloch theorem and describe these AFM systems as they would have only one magnetic site in the unit cell. The validity of this approximation depends on the system: while the parameters $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be comparable to $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the case of CuF2, they appear to be much smaller than $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the case of VF4 (see Fig. 9).

### v2

@@ Acknowledgement
Appendix A Elimination of weakly ferromagnetic components of the spin-orbit coupling
@@
Then, let us consider the unitary transformation of the sublattice $2$, $\hat{U}=e^{i\varphi\boldsymbol{n}\cdot\hat{\boldsymbol{\sigma}}}=\cos\frac{\varphi}{2}\hat{\mathbb{1}}+i\boldsymbol{n}\cdot\hat{\boldsymbol{\sigma}}\sin\frac{\varphi}{2}$, such that $\left(t_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}\right)\hat{U}\approx\tilde{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\mathbb{1}}+it^{y}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\hat{\sigma}_{y}$. In principle, by using such transformation, one can completely eliminate the SO term $i\boldsymbol{t}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\cdot\hat{\boldsymbol{\sigma}}$ separately in each bond [29, 50]. However, this cannot be done simultaneously for all the bonds, unless the parameters of the SO coupling are the same in all these bonds, as the $x$ and $z$ components of the SO coupling responsible for the weak ferromagnetism, for which $t^{x}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv t_{x}$ and $t^{z}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}\equiv t_{z}$. Nevertheless, one can try to eliminate these $t_{x}$ and $t_{z}$.
Assuming $\boldsymbol{n}=\frac{(t_{x},0,t_{z})}{\sqrt{t_{x}^{2}+t_{z}^{2}}}$, which does not depend on the alternating $y$-component, one can find the following equations for $\varphi$:
@@ Appendix B Electronic structure and minimal model for VF4 and CuF2
The transfer integrals operating within the same sublattice were additionally averaged over two sublattices. Namely, using the transfer integrals in the sublattices $I$ and $II$, one can define $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{1}{2}(t^{I}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}+t^{II}_{\boldsymbol{R},\boldsymbol{R}^{\prime}})$ (the regular contribution) and $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}=\frac{1}{2}(t^{I}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}-t^{II}_{\boldsymbol{R},\boldsymbol{R}^{\prime}})$ (the altermagnetic contribution). The approximation consists in neglecting $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$. This artificially suppress the spin-splitting of the AFM bands, but allows us to employ the generalized Bloch theorem and describe these AFM systems as they would have only one magnetic site in cell. The validity of this approximation depends on the system: while the parameters $\delta t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ can be comparable to $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the case of CuF2, they appear to be much smaller than $t^{\phantom{I}}_{\boldsymbol{R},\boldsymbol{R}^{\prime}}$ in the case of VF4 (see Fig. 10).